The perimeter of a rectangle is 460 feet. The length of the rectangle is 40 feet less than three times the width of the rectangle. What are the dimensions of the rectangle?
Answer by ReadingBoosters(3246)

p = 460 = 2w+2l
l = 3w-40
460=2w+2(3w-40)
460 = 2w + 6w - 80
460+80 = 8w
540 = 8w
w = 67.5
l = 3(67.5)-40 = 162.5